# Strange quark matter as dark matter: 40 years later, a reappraisal

### v1

@@ Abstract
Forty years ago, Witten suggested that dark matter could be composed of macroscopic clusters of strange quark matter. This idea was very popular for several years, but it dropped out of fashion once lattice QCD calculations indicated that the confinement/deconfinement transition, at small baryonic chemical potential, is not first order, which seemed to be a crucial requirement in order to produce large clusters of quarks. Here we revisit both the conditions under which strangelets can be produced in the Early Universe and the many phenomenological implications of their existence. Most of the paper discusses the limits on their mass distribution and a possible and simple scheme for their production. Finally, we discuss the most promising techniques to detect this type of objects.
pubyear: 2024pagerange: Strange quark matter as dark matter: 40 years later, a reappraisal12
@@ 1 Introduction
The nature of dark matter remains elusive. One possibility is that dark matter is macroscopic, made of smaller components that can be described within the standard model (Burdin et al., 2015; Jacobs et al., 2015). About forty years ago, it was suggested that dark matter could be made of strangelets, clusters of up, down, and strange quarks (Witten, 1984) formed at the time of the QCD phase transition in the Early Universe. The important point is that dark matter made of strangelets can be effectively weakly interacting because it is extremely massive. This possibility is based on the Bodmer-Witten (BW) hypothesis, stating that clusters of strange quark matter (SQM) can be more stable than the most stable atomic nucleus, i.e. 56Fe. The reason why huge amounts of Fe surround us, while we struggle to find any trace of SQM, is that SQM needs a finite and not-too-small fraction of strange quarks in order to be stable. Since multiple, simultaneous weak decays are enormously suppressed, it is clear that the large amount of strangeness needed to satisfy the BW hypothesis can be obtained only under very specific conditions. To date, the only situations in which those conditions can be satisfied are the Early Universe and the core of massive stellar objects.
@@
What is the theoretical status of the EoSs that can support the BW hypothesis? While it has been known for many years that MIT-bag-like models can satisfy the requests at the basis of that hypothesis (Farhi & Jaffe, 1984), it has been shown that NJL-like models are incompatible with it (Klahn & Fischer, 2015). At the moment, it has been demonstrated that the BW hypothesis can also be realized in the color-dielectric model (Drago & Lavagno, 2001; Alberico et al., 2002; Dondi et al., 2017) and, more generally, in models in which quark masses have a strong density dependence related to confinement. The BW hypothesis can also be realized in models based on perturbative expansions (Kurkela et al., 2010).
In 1984, when Witten wrote his seminal paper, it was considered very plausible that the primordial confinement/deconfinement phase transition was first order. Lattice QCD calculations have later investigated the order of the phase transition in the regime of small baryonic and leptonic chemical potentials, and it has been found that it corresponds to a smooth crossover (Aoki et al., 2006). Recently it has been shown that, if large leptonic asymmetries are taken into account (which is not excluded by cosmology), the transition can indeed be first order (Gao & Oldengott, 2022). It is not yet clear if this will open again the possibility of forming strangelets in the Early Universe following the scheme outlined by Witten. Finally, an important consequence of a first order phase transition in the Early Universe is the production of gravitational waves (Witten, 1984) which could contribute to the signal recently detected by the NANOGrav collaboration (Afzal et al., 2023).
@@ 2 Observational constraints
Strangelets belong to the macros category, which has to respect several observational constraints from different sources (Jacobs et al., 2015; Burdin et al., 2015). The mass ranges not subject to limits for nuclear density monochromatic macros in the present Universe are $(55-10^{3})$ g, $(5\times 10^{4}-10^{8})$ g and $(10^{10}-10^{18})$ g (Singh Sidhu & Starkman, 2020; Burdin et al., 2015). In the case of strangelets, other constraints must also be taken into account, because they can interact with ordinary matter, dramatically altering the astrophysical evolution of stars. In particular, Madsen (1988) rules out the two lowest mass windows, leaving open (at least for monochromatic strangelets) only the range $(10^{10}-10^{18})$ g.
@@ 3.1 Evaporation
The mechanism of evaporation primarily involves the emission of nucleons from the strangelet. This process is regulated by various mechanisms, described in the following as outlined in Alcock & Farhi (1985), in particular: the thermal balance of the strangelet which cools down by the evaporation process itself and needs to be kept warm by neutrinos; the net-rate of emission of nucleons by an isolated strangelet; the increase in the re-absorption rate due to the presence of evaporated nucleons close to the surface of the strangelet; the opacity of the evaporated material to neutrinos. In more details:
@@
    Finally, given the relationship between the age of the universe $\tau_{U}$ and its temperature $T_{u}$ in the radiation-dominated era (Alcock & Farhi, 1985):
@@ 3.2 Strangelet distributions
To accurately model the current distribution of dark matter in the galaxy as strangelets, one must consider various factors, including the strangelet size distribution. Initially, we hypothesize a log-normal distribution of sizes and then subject it to an evaporation process. This evaporation, occurring in a range of temperatures from 150 MeV to 1 MeV, results in a transformed distribution that must agree with the constraints specified in Section 2.
@@ 3.2.1 Mathematical framework
It is crucial to establish a mathematical framework for handling the size distributions of strangelets. Here we will follow the scheme deveoped in Bucciantini et al. (2022). Given an initial distribution $P(A^{\prime})$, representing the number of strangelets with a baryon number $A^{\prime}$, the distribution after evaporation, denoted as $Q(A)$, is given by:
@@ 3.2.3 Galactic distribution
Figure 3: This figure showcases distributions with varying number densities and mass densities of strangelets. The represented curves, in red, green, yellow, grey, purple, and blue, correspond to distributions containing approximately $10^{34}$, $10^{33}$, $10^{32}$, $10^{31}$, $10^{30}$, and $10^{29}$ strangelets, respectively. Notably each curve, despite having different peak densities, has the same total mass. Furthermore, each distribution agrees with the femtolensing observational constraints and other relevant ones as detailed in Singh Sidhu & Starkman (2020) and Burdin et al. (2015). The distributions are normalized to the estimated dark matter mass in the Milky Way (Nesti & Salucci, 2013).
@@ 3.3 Rephrasing Witten with a modern language
As already discussed, some of the hypotheses assumed in Witten (1984) and Farhi & Jaffe (1984) are probably not compatible with what is now known about microphysics. In particular the following questions were not discussed, mostly because forty years ago these problems were still in a very primordial stage or completely unknown: chiral dynamics across the critical temperature; compatibility of the value of the bag pressure with the results of experiments of heavy ion collisions; role played by the formation of diquark condensates. In this section we discuss a possible path, incorporating state-of-the-art knowledge of quark dynamics which, although speculative, could lead to the formation of strangelets in a way which is phenomenologically not too different from the one based on the papers of Witten (1984) and Farhi & Jaffe (1984). It is indeed interesting to remark that the distribution of the size of the strangelets, obtained in the previous section by taking into account all observational limits, peaks at sizes very close to the one estimated in the paper by Witten (1984). It is therefore important to investigate whether a cosmological path, incorporating up-to-date results of microphysics, can still support that result.
We propose a scheme based on the following stages:
@@
  • lattice QCD simulations show that, at temperatures larger than the critical one, strong correlations between quarks can take place (Bellwied et al., 2021). If the cosmological trajectory is not too far from the critical end point (Gao & Oldengott, 2022), these correlations can correspond to the formation of pre-clusters, as discussed in Shuryak & Torres-Rincon (2020). Notice that, at this stage, chiral symmetry is still restored and therefore the masses of up, down and strange quarks are still small;
@@
  • while the transition is not assumed to be first order, we imagine that the decrease of the temperature can lead to an increase of the correlations and to the gradual development of binding energy in the cluster. Once the (pre-)cluster is bound, its total energy can be described by a semi-empirical mass formula, including also a surface term. While this step is still speculative, the formation of nuclei in heavy ion collisions (Adam et al., 2016; Braun-Munzinger & Dönigus, 2019) supports this hypothesis since it shows that the formation of bound clusters can take place even if a first order phase transition is not observed;
  • the size of clusters can dramatically increase, by a process of coalescence totally similar to the one discussed by Witten (1984). In this way, the initial distribution of the size of the strangelets should peak in the same range indicated by Witten (1984);
  • when the temperature further decreases, chiral symmetry breaks. The quark masses increase, including the mass of the strange quark, and the cluster becomes meta-stable. The presence of a surface energy implies that the clusters are not immediately destroyed. The process of evaporation is also suppressed because neutrinos (crucial to evaporation, as discussed in the previous section) are absorbed by the layer of evaporated hadrons and they are therefore unable to keep the cluster hot (Alcock & Farhi, 1985);
  • since up, down and strange quarks are already present in the cluster, nothing stops the cluster from increasing its density in order to minimize its energy by forming a phase in which quarks are gapped. At that point, the clusters are totally stable and they follow the partial evaporation path described in the previous section. Notice that once clusters form, large enough to be described through hydrodynamics, the process of conversion from normal quark matter to gapped quark matter is very rapid, due to the onset of hydrodynamical instabilities (Drago et al., 2007). This mechanism of stabilizing clusters of SQM has been addressed in several papers, including the ones discussing the evolution of an entire strange quark star Alford & Reddy (2003); Drago et al. (2004).
@@
A macroscopic isolated strangelet develops a Coulomb barrier, similarly to what happens in the case of a strange quark star (Alcock et al., 1986). The height of the barrier depends on the temperature (Kettner et al., 1995) and in Fig. 5 we show this dependence. At low temperatures ($T\lesssim 1\mathrm{MeV}$) the barrier is large enough to stop the absorption of protons. Instead, for $T\sim$ a few ten MeV (the range relevant for evaporation), the barrier is already low compared to the temperature and therefore it can be neglected when discussing proton re-absorption (see subsection 3.1). Finally, for $T\gtrsim 100$ MeV at which clusterization still takes place, the Coulomb barrier has not yet formed and the value indicated in Fig. 5 corresponds just to an upper limit.
Of course the scheme discussed above is speculative and detailed analyses will be needed in order to clarify all the steps. For instance, it is crucial to understand the evolution of the (pre-)clusters with the temperature. Also, the cosmic trajectory has, up to now, been computed assuming a small contribution of multistrange particles (Letessier & Rafelski, 2002). It will be interesting to investigate if some of the models used to describe the formation of nuclei at high temperature, as observed in heavy ion collision experiments, can be adapted to the cosmological scenario. Another important point has to do with the formation of gapped quark matter. We do not know if the gapping will directly produce Color-Flavor-Locked quark matter Alford et al. (2008), or if it will form, at least as an intermediate step, a crystalline phase as the one described in Anglani et al. (2014). Notice that the knowledge of which gapping pattern will actually take place is important also for the phenomenology of dark matter, since Color-Flavor-Locked quark matter is totally neutral, while a crystalline phase needs the presence of electrons to achieve charge neutrality.

### v3

@@ Abstract
Forty years ago Witten suggested that dark matter could be composed of macroscopic clusters of strange quark matter. This idea was very popular for several years, but it dropped out of fashion once lattice QCD calculations indicated that the confinement/deconfinement transition, at small baryonic chemical potential, is not first order, which seemed to be a crucial requirement in order to produce large clusters of quarks. Here we revisit the conditions under which strangelets can be produced in the Early Universe. We discuss the impact of an instability in the hadronic phase separating a low density, positive-strange-charge phase from a high-density phase with a negative strange charge. This second phase can rapidly stabilize by forming color-superconducting gaps. The strangelets then undergo partial evaporation. In this way, we obtain distributions of their sizes in agreement with the observational constraints and we discuss the many astrophysical and cosmological implications of these objects. Finally, we examine the most promising techniques to detect this type of strangelets. We also show that strangelets can exist with masses $\lesssim 10^{17}\mathrm{g}$, while primordial black holes are ruled out in that mass range, allowing us to distinguish between these two dark matter candidates.
keywords:
cosmology: dark matter - cosmology: early Universe - dense matter.
pubyear: 2024pagerange: Strange quark matter as dark matter: 40 years later, a reappraisal13
@@ 1 Introduction
The nature of dark matter remains elusive, also for its derived distribution in galaxies (Salucci, 2019). One possibility is that dark matter is macroscopic, made of smaller components that can be described within the standard model (Burdin et al., 2015; Jacobs et al., 2015). About forty years ago, it was suggested that dark matter could be made of strangelets, clusters of up, down, and strange quarks (Witten, 1984) formed at the time of the QCD phase transition in the Early Universe. The important point is that dark matter made of strangelets can be effectively weakly interacting because it is extremely massive. This possibility is based on the Bodmer-Witten (BW) hypothesis (Bodmer, 1971; Witten, 1984), stating that clusters of strange quark matter (SQM) can be more stable than the most stable atomic nucleus, i.e. 56Fe. The reason why huge amounts of Fe surround us, while we struggle to find any trace of SQM, is that SQM needs a finite and not-too-small fraction of strange quarks in order to be stable. Since multiple, simultaneous weak decays are enormously suppressed, it is clear that the large amount of strangeness needed to satisfy the BW hypothesis can be obtained only under very specific conditions. To date, the only situations in which those conditions can be satisfied are the Early Universe and the core of massive stellar objects.
It is important to note that the validity of the BW hypothesis does not imply that dark matter is made of SQM: it is possible that the hadronization process in the Early Universe does not allow the production of the large clusters of SQM that would constitute dark matter (the possibility that dark matter is made of small clusters of SQM is probably ruled out by observations, as we will discuss in the following).
@@
What is the theoretical status of the EoSs that can support the BW hypothesis? While it has been known for many years that MIT-bag-like models can satisfy the requests at the basis of that hypothesis (Farhi & Jaffe, 1984), it has been shown that NJL-like models are incompatible with it (Klahn & Fischer, 2015). At the moment, it has been demonstrated that the BW hypothesis can also be realized in the color-dielectric model (Drago & Lavagno, 2001; Alberico et al., 2002; Dondi et al., 2017) and, more generally, in models in which quark masses have a strong density dependence related to confinement. A crucial role in these analyses can be played by the gluon dynamics in the infrared regime (Cyrol et al., 2018; Huber, 2020). In particular, the gluon condensate can act as an effective gluon mass, regularizing the infrared behavior (Adler & Piran, 1984; Horak et al., 2022). The BW hypothesis can also be realized in models based on perturbative expansions (Kurkela et al., 2010).
In 1984, when Witten wrote his seminal paper, it was considered very plausible that the primordial confinement/deconfinement phase transition was first order. Lattice QCD calculations have later investigated the order of the phase transition in the regime of small baryonic and leptonic chemical potentials, and it has been found that it corresponds to a smooth crossover (Aoki et al., 2006).It is still expected that a critical endpoint exists, in the chemical potential - temperature plane, and that to the right of that point the transition is a first order Stephanov (2004). Recently, it has been shown that, if large leptonic asymmetries are taken into account (which is not excluded by cosmology (Froustey & Pitrou, 2024)), the trajectory followed by the Early Universe can indeed pass through a first order transition (Gao & Oldengott, 2022). It is not yet clear if this will open again the possibility of forming strangelets in the Early Universe following the scheme outlined by Witten. Finally, an important consequence of a first order phase transition in the Early Universe is the production of gravitational waves (Witten, 1984) which could contribute to the signal recently detected by the NANOGrav collaboration (Afzal et al., 2023).
@@ 2 Observational constraints
Strangelets belong to the macros category, which has to respect several observational constraints from different sources (Jacobs et al., 2015; Burdin et al., 2015). The mass ranges not subject to limits for nuclear density monochromatic macros in the present Universe are $(55-10^{3})$ g, $(5\times 10^{4}-10^{8})$ g and $(10^{10}-10^{18})$ g (Singh Sidhu & Starkman, 2020; Burdin et al., 2015). In the case of strangelets  [^1] Other types of macros composed of quarks/antiquarks have been proposed, for instance the ones associated with the collapse of domain walls Bai et al. (2023) and Liang & Zhitnitsky (2016)., other constraints must also be taken into account, because they can interact with ordinary matter, dramatically altering the astrophysical evolution of stars. In particular, Madsen (1988) rules out the two lowest mass windows, leaving open (at least for monochromatic strangelets) only the range $(10^{10}-10^{18})$ g.
@@ 2.6 Proto-neutron stars
2.7 Comparison with primordial black-holes
Primordial black holes (PBHs) have long been considered as potential dark matter candidates (Carr & Hawking, 1974; Carr, 1975; Carr et al., 2021), and many of the observational limits that restrict strange quark matter as a dark matter component also apply to PBHs. Constraints from gravitational microlensing surveys such as MACHO, EROS and OGLR (Alcock et al., 2000; Tisserand et al., 2007; Niikura et al., 2019), and from femtolensing (Barnacka et al., 2012) shape the allowed mass ranges of both PBHs and strange quark nuggets.
Despite these parallels, one crucial difference is due to the possible evaporation of PBHs. Black holes with masses smaller than $\sim 10^{17}\,\mathrm{g}$ are effectively ruled out as significant dark matter constituents due to Hawking radiation (Hawking, 1974, 1975). Such low-mass PBHs would have evaporated well before the present epoch, depleting their abundance. In contrast strangelets, do not undergo a Hawking-like evaporation process and therefore they are not subject to this constraint.
@@ 3.1 Evaporation
The mechanism of evaporation primarily involves the emission of nucleons from the strangelet. This process is regulated by various mechanisms, as outlined in Alcock & Farhi (1985), in particular: the thermal balance of the strangelet which cools down by the evaporation process itself and needs to be kept warm by neutrinos; the net-rate of emission of nucleons by an isolated strangelet; the increase in the re-absorption rate due to the presence of evaporated nucleons close to the surface of the strangelet; the opacity of the evaporated material to neutrinos. More in details:
@@
    Finally, given the relation between the age of the universe $\tau_{U}$ and its temperature $T_{u}$ in the radiation-dominated era (Alcock & Farhi, 1985):
@@ 3.2 Strangelet distributions
To accurately model the current distribution of dark matter in the galaxy as strangelets, one must consider various factors, including the strangelet number distribution. Initially, we hypothesize a log-normal distribution of sizes and then subject it to an evaporation process. This evaporation, occurring in a range of temperatures from 150 MeV to 1 MeV, results in a transformed distribution that must agree with the constraints specified in Section 2.
@@ 3.2.1 Mathematical framework
It is crucial to establish a mathematical framework for handling the number distributions of strangelets. Here we will follow the scheme developed in Bucciantini et al. (2022). Given an initial distribution $P(A^{\prime})$, representing the number of strangelets with a baryon number $A^{\prime}$, the distribution after evaporation, denoted as $Q(A)$, is given by:
@@ 3.2.3 Galactic distribution
Figure 3: Number distributions (solid) and mass distributions (dashed) of strangelets. The curves in red, green, yellow, gray, purple, and blue, correspond to distributions containing approximately $10^{34}$, $10^{33}$, $10^{32}$, $10^{31}$, $10^{30}$, and $10^{29}$ strangelets, respectively. The distributions are consistent with the observational constraints detailed in Section 2 and have been normalized to the estimated dark matter mass in the Milky Way (Nesti & Salucci, 2013).
@@ 3.3 Rephrasing Witten with a modern language
As already discussed, some of the hypotheses assumed in Witten (1984) and Farhi & Jaffe (1984) are probably not compatible with what is now known about microphysics. In particular the following questions were not discussed, mostly because forty years ago these problems were still in a very primordial stage or completely unknown: chiral dynamics across the critical temperature; compatibility of the value of the bag pressure with the results of experiments of heavy ion collisions; role played by the formation of diquark condensates. In this section we discuss a possible path which, although speculative, incorporates state-of-the-art knowledge of quark dynamics and could lead to the formation of strangelets in a way which is phenomenologically not too different from the one based on the papers of Witten (1984) and Farhi & Jaffe (1984). It is indeed interesting to remark that the distribution of the sizes of the strangelets, obtained in the previous section by taking into account all observational limits, peaks at sizes very close to the one estimated in the paper by Witten (1984). It is therefore important to investigate whether a cosmological path, incorporating up-to-date results of microphysics, can still support that result.
We propose a scheme based on the following stages, see Fig. 5:
  1. 1.
@@
    lattice QCD simulations show that, at temperatures larger than the critical one, strong correlations between quarks can already take place (Bellwied et al., 2021). Notice that, at this stage, chiral symmetry is still restored and therefore the masses of up, down and strange quarks are still small;
  2. 2.
@@
    while the transition is not assumed to be first order, we imagine that the decrease of the temperature can lead to an increase of the correlations and to the gradual development of binding energy in the cluster, with a mechanism similar to quark-coalescence (Song & Coci, 2022; Aaij et al., 2024). Once the cluster is bound, its total energy can be described by a semi-empirical mass formula, including also a surface term. If the cosmological trajectory is not too far from the critical end point (Gao & Oldengott, 2022), these correlations can develop more easily;
  3. 3.
    the size of the clusters, initially of the order of few femtometers, can dramatically increase, by a process of cluster coalescence totally similar to the one discussed by Witten (1984). This process could start with the fusion of clusters of three quarks, maybe by using clusters having the quantum numbers of the sexaquark as the doorway to the formation of much larger clusters (Farrar, 2022; Shahrbaf et al., 2022). This process should be mildly exothermic and therefore the temperature should remain roughly constant. A potential problem with this path is that, unless the new phase is stable under weak interactions, it can decay on a timescale shorter than the one needed to form large clusters. Another possible path leading to the formation of larger clusters is based on the existence of chemical-mechanical instabilities, in the hadronic phase, at temperatures close to the critical one. This possibility has been discussed in Lavagno & Pigato (2022), where it has been shown that a phase separation can take place, between a low-density, positive-strange-charge phase and a higher density phase with a negative strange charge. This second phase can form, on the timescale of the strong interaction, as soon as the system starts hadronizing, it has a finite content of hyperons and it can constitute the bridge to the formation of diquark matter described in the next stage. The exact location of this instability still needs to be worked out (and is model dependent), but it could start at a chemical potential smaller than that of the QCD critical point (see red segment in Fig. 5). This is not surprising, since it is known, also from lattice calculations, that phase transitions in the hadronic phase can take place at very low densities, for instance for isospin asymmetric matter (Brandt et al., 2018; Vovchenko et al., 2021);
  4. 4.
    even if in the previous stage large clusters can already form, they can be meta-stable and therefore rapidly evaporate. This problem could be solved by the formation of color superconducting energy gaps (Alford et al., 2008; Lugones & Horvath, 2004), but the maximum temperature at which these gaps can form is often estimated to be significantly smaller than the deconfinement temperature. The existence of a rather large gap between these two temperatures would lead to a rapid evaporation of the clusters. This problem was noticed by Zhitnitsky (2003) and led to the suggestion that the clusters could be stabilized by the pressure of axion domain walls. We propose instead that the value of the superconducting gap, and therefore of the associated critical temperature, is significantly larger than the one normally discussed in the literature. Notice that most of the estimates present in the literature are based on the NJL model (Ruester et al., 2006), but a similar calculation based on the PNJL model suggests that the diquarks can form at temperatures not much smaller than the critical one (Roessner et al., 2007). Also a preliminary calculation within the chromodielectric model suggests significantly larger gaps, whose value increases at low densities (de Carvalho et al., 2010). At variance with NJL, PNJL and the chromodielectric are confining models and this can play a crucial role in order to obtain absolutely stable strange quark matter (Drago & Lavagno, 2001; Dondi et al., 2017). Notice that since up, down and strange quarks are already present in the cluster, nothing stops the cluster from forming a phase in which quarks are in a CFL phase. This process can take place only if the clusters are larger than several femtometers, otherwise the gap formation is suppressed (Amore et al., 2002). At that point, the clusters are totally stable and they follow the partial evaporation path described in the previous section. This mechanism of stabilizing clusters of SQM has been addressed in several papers, including the ones discussing the evolution of an entire strange quark star (Alford & Reddy, 2003; Drago et al., 2004). Stages iii and iv can be summarized by saying that we assume that a phase transition takes place, associated with a mechanical instability, which leads to the formation of a new phase having a large baryon density and a finite and negative strange charge. This new phase can rapidly transform into a CFL phase, which can be particularly easy if there is no phase transition between hypernuclear matter and CFL quark matter (Schäfer & Wilczek, 1999; Alford et al., 1999, 2008). In this way we have re-introduced a phase transition with a large density jump, similar to the one discussed by Witten (1984), but taking place at temperatures slightly smaller than the critical one in the already hadronized matter. Also in the scenario discussed here, gravitational waves could be generated during this process (Witten, 1984; EPTA Collaboration et al., 2024).
  5. 5.
    The process of evaporation is initially suppressed because neutrinos (crucial to evaporation, as discussed in the previous section) are absorbed by the layer of evaporated hadrons and they are therefore unable to keep the cluster hot (Alcock & Farhi, 1985). On the other hand, evaporation is crucial to fix the right balance between ordinary matter (the evaporated part) and dark matter (the remaining large clusters).
Figure 5: Cosmological path of quark clusters, following the stages outlined in subsection 3.3. The phase separation is schematic and is loosely based on Roessner et al. (2007). The solid line indicates a first-order phase transition between the chiral symmetry broken phase and the diquark phase; the dashed line a crossover and the dot-dashed a second order. This last transition, separating the diquark phase from the quark gluon plasma (QGP), can be much more complicated then the schematic one shown in the figure, and its order is strongly model dependent. The stages include: (1) initial quark correlations above the critical temperature, (2) growth of binding energies and quark-coalescence, (3) formation of large clusters, (4) stabilization of the clusters by color superconducting gaps and (5) partial evaporation. The red segment indicates the instability discussed in Lavagno & Pigato (2022). If that instability exists and produces hadronic matter with a finite and negative strange charge it is possible that the transition between hadronic matter and CFL phase is totally continuous, see text.
The scheme outlined above is compatible with the evaporation process described in the previous section. In particular, the formation of diquark condensates at large temperatures is crucial to avoid the rapid evaporation of the clusters.
@@
Up to now we have not discussed the role possibly played by the Coulomb interaction during coalescence. A macroscopic isolated strangelet develops a Coulomb barrier, similarly to what happens in the case of a strange quark star (Alcock et al., 1986). The height of the barrier depends on the temperature (Kettner et al., 1995) and in Fig. 6 we show this dependence. At low temperatures ($T\lesssim 1$ MeV) the barrier is large enough to stop the absorption of protons. Instead, for $T\sim$ a few ten MeV (the range relevant for evaporation), the barrier is already low compared to the temperature and therefore it can be neglected when discussing proton re-absorption (see subsection 3.1). Finally, for $T\gtrsim 100$ MeV at which clusterization still takes place, the Coulomb barrier has not yet formed and the value indicated in Fig. 6 corresponds just to an upper limit.
Of course the scheme discussed above is speculative even though the formation of nuclei in heavy ion collisions (Adam et al., 2016; Braun-Munzinger & Dönigus, 2019) somehow supports this scheme, since it shows that the formation of bound clusters can take place even if a first order phase transition is not observed. Detailed analyses will be needed in order to clarify all the steps. For instance, it is crucial to understand the evolution of the clusters with the temperature. Also, the cosmic trajectory has, up to now, been computed assuming a small contribution of multistrange particles (Letessier & Rafelski, 2002). It will be interesting to investigate if some of the models used to describe the formation of nuclei at high temperature, as observed in heavy ion collision experiments, can be adapted to the cosmological scenario. Another important point has to do with the formation of gapped quark matter. We do not know if the gapping will directly produce Color-Flavor-Locked quark matter (Alford et al., 2008), or if it will form, at least as an intermediate step, a crystalline phase as the one described in Anglani et al. (2014). Notice that the knowledge of which gapping pattern will actually take place is important also for the phenomenology of dark matter, since Color-Flavor-Locked quark matter is totally neutral, while a crystalline phase needs the presence of electrons to achieve charge neutrality.